# Non-Adaptive Multi-Stage Algorithm and Bounds for Group Testing with Prior Statistics

### v4

@@ III-B Analytical Results
Let $G$ denote the number of non-defective items that were hidden and not detected in DND (i.e., the number of false positive items). We upper bound the probability of error in DND, by the sum of the average probability of error $P_{e,a}^{(DND)}$ and a concentration term as follows:
$P_{e,u}^{(DND)}\triangleq P_{e,a}^{(DND)}+P\left(G-\mathbb{E}[G]>g\right),$
for some $g>0$.
Although the variation of LVA in Stage 2 applies a maximum likelihood decision, it’s performance in the proposed algorithm, may be worse than a brute force MAP’s performance, if the number of occlusions in the test matrix exceeds the threshold $\gamma K$. For this reason we are interested in $g=\gamma K-\left(K+\left(N-K\right)P_{e,a}^{(DND)}\right)$.
Recall that the probability of each definitely defective item to be hidden is given by the probability of error in DND, $P_{e}^{(DND)}$, which depends only on the matrix design. Thus, in the case of Bernoulli encoder, the $N-K$ random variables which represent occlusions of non-defective items are independent. It follows that $G\sim Bin\left(N-K,P_{e}^{(DND)}\right)$, and we can apply the one-sided Chebyshev’s inequality [52] to bound the concentration term:
$P\left(G-\mathbb{E}\left[G\right]>g\right)\leq\\ \left(1+\frac{g^{2}}{\left(N-K\right)P_{e}^{(DND)}\left(1-P_{e}^{(DND)}\right)}\right)^{-1}$
For a sufficiently large $T$, $P_{e}^{(DND)}\rightarrow P_{e,a}^{(DND)}$, which concludes the proof. ∎
@@
Consider a group test with a Bernoulli testing matrix with $p=\ln{2}/K$, and $T$ tests as $K\rightarrow\infty$. The expected number of defective items that DD successfully detects is bounded by
$\mathbb{E}\left[\left|\mathcal{P}^{\left(DD\right)}\right|\right]\geq K\left(1-N^{-\frac{\alpha}{2}\left(1-\ln{2}/K\right)^{N\cdot P_{e,a}^{(DND)}}}\right).$ (2)
@@
The detection of the $i$-th defective item ($i\in\mathcal{K}$) in the DD algorithm fails under two conditions: either it does not participate in any test, or it participates but at least one other potentially defective item is also participating. Let $A_{i}$ represent the event where at least one potentially defective item, excluding item $i$, participates in some test. Consequently, $P\left(A_{i}\right)=1-\left(1-p\right)^{\left|\mathcal{P}^{(S_{1,1})}\right|-1}$. The probability of not identifying item $i$ as definitively defective in a given test $t\in\{1,...,T\}$ is given by $1-p+pP\left(A_{i}\right)$. So the probability of not detecting a defective item in all tests is given by
$P_{e}^{(DD)}$ $=\left[1-p+p\left(1-\left(1-p\right)^{\left|\mathcal{P}^{(S_{1,1})}\right|-1}\right)\right]^{T}$
$=\left[1-p\left(1-p\right)^{\left|\mathcal{P}^{(S_{1,1})}\right|-1}\right]^{T}.$
By substituting (1) we have:
$\mathbb{E}\left[P_{e}^{(DD)}\right]$
$\leq\left(1-p\left(1-p\right)^{K+\left(N-K\right)P_{e,a}^{(DND)}-1}\right)^{T}$
$\leq\left(1-p\left(1-p\right)^{K+P_{e,a}^{(DND)}N-1}\right)^{T}$
=$\underset{(a)}{=}\left(1-\frac{\ln{2}}{K}\left(1-\frac{\ln{2}}{K}\right)^{K-1}\left(1-\frac{\ln{2}}{K}\right)^{P_{e,a}^{(DND)}N}\right)^{T}$
$\underset{(b)}{\leq}\left(1-\frac{\ln{2}}{K}e^{-\ln{2}}\left(1-\frac{\ln{2}}{K}\right)^{P_{e,a}^{(DND)}N}\right)^{\alpha K\log_{2}{N}}$
$\underset{(c)}{\leq}exp\left(-\alpha e^{-\ln{2}}\ln{2}\left(1-\frac{\ln{2}}{K}\right)^{P_{e,a}^{(DND)}N}\log_{2}{N}\right)$
$=N^{-\frac{\alpha}{2}\left(1-\frac{\ln{2}}{K}\right)^{P_{e,a}^{(DND)}N}},$
where in $(a)$ we substitute $p=\ln{2}/K$, in $(b)$ we substitute $T=\alpha K\log_{2}{N}$ and used the fact that $e^{-\ln 2}\leq\left(1-\ln{2}/n\right)^{n-1}$ for any integer $n>0$. $(c)$ follows since $e^{-x}\geq\left(1-x/n\right)^{n}$ for any integer $n>0$ and $x>0$.
Subsequently, the probability of success satisfies
$\mathbb{E}\left[P_{s}^{(DD)}\right]=1-\mathbb{E}\left[P_{e}^{(DD)}\right]\geq 1-N^{-\frac{1}{2}\alpha\left(1-\frac{\ln{2}}{K}\right)^{P_{e,a}^{(DND)}N}}.$
Thus, an upper bound on the expected number of defective items that the DD algorithm successfully detects can be established as follows:
$\mathbb{E}\left[\left|\mathcal{P}^{\left(DD\right)}\right|\right]=\mathbb{E}\left[P_{s}^{(DD)}\right]K\geq\left(1-N^{-\frac{1}{2}\alpha\left(1-\frac{\ln{2}}{K}\right)^{P_{e,a}^{(DND)}N}}\right)K.$
This completed the theorem proof. ∎
The proposition below provides a sufficient condition on $\gamma$, such that, MSGT achieves performance that outperforms ML. This can be assumed since LVA used for this step is an optimal MAP estimator [32].
Proposition 1.
Let $\eta$ denote the average success probability of ML decoder. For GT with correlated prior information and $\gamma\geq 1$, MSGT achieve an average success probability equal or greater than $\eta$, if
$T\geq\log{\eta}+\log{\binom{N}{K}}-\log{\binom{N-K}{K(\gamma-1)}}.$ (3)
@@
We follow the lines of the success probability analysis in [45, Theorem 1]. Consider a list decoder that, given the test outcome $Y$, outputs $\mathcal{L}(Y)$, a list of items estimated as defective, such that $\lvert\mathcal{L}(Y)\lvert\geq K$. Let $N(Y)$ denote the number of possible defective sets for which, given $Y$, the list decoding is successful: $N(Y)\triangleq\sum_{\mathcal{K}}\mathbb{1}_{\{\mathcal{L}(Y)\cap\mathcal{K}=K\}}$. Assuming no prior information is given, $N(Y)$ is bounded by the counting bound
$N(Y)\leq\binom{N-K}{L-K}.$ (4)
The success probability of the decoder can be written as
$P_{s}=\frac{1}{\binom{N}{K}}\sum_{Y}N(Y)=\frac{2^{T}}{\binom{N}{K}}N(Y).$ (5)
By substituting (4) into (5), taking the logarithm and rearranging, we have that
$T\geq\log{P_{s}}+\log{\binom{N}{K}}-\log{\binom{N-K}{L-K}}.$
It is important to note that unlike [45, Theorem 1], we halt the success probability analysis here without deriving a converse, and instead focus on the connection between $T$ and $\gamma$. This analysis is valid for any list decoder, particularly for an ML decoder [17, 53], allowing us to set $P_{s}=\eta$. Since MSGT uses MAP, the non-uniform and correlated prior information allows us to achieve better recovery performance for the same $T$ and $L$ compared to ML [^2] In Section IV-C, we provide sufficiency bound for any MAP decoder with an arbitrary underlying correlation between infected items that outperform the ML sufficiency bound and illustrate it for practical scenarios in Section V.. Moreover, since for any $\gamma\geq 1$, $\gamma K$ is the maximum list size of estimated defective items we allow in MSGT. By substituting $L=\gamma K$, we establish a lower bound for $\gamma$ that guarantees MSGT’s success probability remains at least equal to that of the ML estimator. ∎
One of the key features of the proposed MSGT algorithm is its low and feasible complexity in practical regimes compared to ML or MAP-based GT decoders. Both ML and MAP involve exhaustive searches, resulting in a complexity of $\mathcal{O}(\binom{N}{K}KN\log_{2}N)$ operations [17]. The theorem below and the subsequent remark characterize the computational complexity of MSGT.
Theorem 4.
Consider a group test for a population of $N$ items, of which $K$ are defective and a Bernoulli testing matrix. The computational complexity of the MSGT algorithm is bounded by $\mathcal{O}\left(L\gamma^{K}KN\log_{2}N\right)$ operations.
Proof.
@@
We begin by analyzing the complexity of each step of the proposed MSGT solution given in Algorithm 1, and finally, we sum everything up to determine the total complexity.
@@
In Stage 1, the complexity of DND is $\mathcal{O}\left(KN\log_{2}{N}\right)$ as analyzed in [54, Remark 6]. Then, for each positive entry of the test result vector $Y^{T}$, the DD algorithm counts the number of possibly defective items that participate in the corresponding pool test. That requires $K\left|\mathcal{P}^{(S_{1,1})}\right|\log_{2}{N}$ computations, and for simplification, we bound it by the DND complexity, i.e., $\mathcal{O}\left(KN\log_{2}N\right)$.
In Stage 2, parallel LVA requires $L$ times more computations than the VA [32]. The VA calculates all the possible transition probabilities for each step in the sequence. In GT, this sequence is the order’s items sequence, where with the suggested algorithm, it is enough to consider only the $\left|\mathcal{P}^{\left(S_{1,2}\right)}\right|$ items as the sequence steps. The optional states are basically either “non-defective” or “defective”, so there are four possible transitions in each step of the trellis proposed herein for GT. Nevertheless, this algorithm can be implemented to leverage additional memory to decide the state of each item based on the preceding $\tau$ items. Consequently, LVA takes $2^{2\tau}L\left|\mathcal{P}^{(S_{1,2})}\right|$ computations. For the average case, we use the expectation bounds from Theorems 1 and 3, such that:
$\mathbb{E}\left[\left|\mathcal{P}^{\left(S_{1}\right)}\right|\right]=\mathbb{E}\left[\left|\mathcal{P}^{(S_{1,1})}\right|\right]-\mathbb{E}\left[\left|\mathcal{P}^{\left(DD\right)}\right|\right]\leq\left(N-K\right)P_{e,a}^{(DND)}+N^{-\frac{\alpha}{2}\left(1-\ln{2}/K\right)^{N\cdot P_{e,a}^{(DND)}}}K.$
Accordingly, the number of the required computations for LVA step in MSGT is bounded by
$2^{2\tau}L\left[\left(N-K\right)P_{e,a}^{(DND)}+N^{-\frac{\alpha}{2}\left(1-\ln{2}/K\right)^{N\cdot P_{e,a}^{(DND)}}}K\right].$ (6)
For the worst-case, we assume that the DD step does not affect the possible detected items set, thus $\left|\mathcal{P}^{\left(S_{1,2}\right)}\right|\leq\left|\mathcal{P}^{(S_{1,1})}\right|$. To further bound the expression, we utilize both the error probability of the worst case as given in Theorem 2, thus $\left|\mathcal{P}^{\left(DND\right)}\right|\leq K+(N-K)P_{e,u}^{(DND)}$. By substituting Theorem 2 it follows that the number of possibly defective items in the worst-case is upper bounded by
$\begin{split}K&+\left(N-K\right)\left(P_{e,a}^{(DND)}+\left(1+\frac{g^{2}}{\left(N-K\right)P_{e,a}^{(DND)}\left(1-P_{e,a}^{(DND)}\right)}\right)^{-1}\right),\end{split}$ (7)
for $g=\gamma K-\left(K+\left(N-K\right)P_{e,a}^{(DND)}\right)$. Now, since the two expressions multiplying $(N-K)$ in (6) and (7) are error probabilities, both of them can be roughly bounded by one. Hence, the number of computations of LVA in the average case and worst-case becomes $\mathcal{O}(LN)$.
@@
In the next step in Stage 2, we filter the LVA results. We sum each sequence $\mathbf{Z_{l}}$ with a complexity of $\mathcal{O}(N)$. If the result is in the range $\left[K,\gamma K\right]$, we extract all the combinations of size $K-\left|\mathcal{P}^{(DD)}\right|$. Thus, this stage is done in at most $\mathcal{O}\left(L\left(N+\binom{\gamma K}{K-\left|\mathcal{P}^{(DD)}\right|}\right)\right)$ computations.
Finally, in the MAP step of Stage 2, the algorithm goes over at most $L$ combinations of size $K-\left|\mathcal{P}^{(DD)}\right|$ out of no more than $\gamma K$ possibly defective items in each sequence. Then, for each combination, the group test is applied. Therefore, the MAP stage requires $\binom{L\gamma K}{K}KN\log_{2}{N}$ computations. Substituting the bounds of the binomial coefficient $\left(\frac{\gamma K}{K}\right)^{K}\leq\binom{\gamma K}{K}\leq\left(\frac{e\gamma K}{K}\right)^{K}$, it follows that the complexity of MAP stage in the proposed MSGT is $\mathcal{O}\left(L\gamma^{K}KN\log_{2}{N}\right)$.
To conclude, the complexity of MSGT is the sum of all the steps, i.e., $\mathcal{O}\big{(}NK\log_{2}{N}+LN+L\binom{\gamma K}{K-\left|\mathcal{P}^{(DD)}\right|}+L\gamma^{K}KN\log_{2}{N}\big{)}$. As $N$ grows, the dominant term is the complexity of the MAP step. Thus, the complexity of the MSGT algorithm is bounded by $\mathcal{O}\left(\gamma^{K}KN\log_{2}{N}\right)$ operations, which completes the proof. ∎
@@
If we skip the LVA step, MSGT converges to the MAP estimator. Thus, the MAP’s complexity is
$\mathcal{O}\left(\displaystyle\binom{\left|\mathcal{P}^{(S_{1,2})}\right|}{K-\left|\mathcal{P}^{(DD)}\right|}KN\log_{2}N\right),$
when the DND and DD are executed as prior steps, and otherwise it is $\mathcal{O}\left(\binom{N}{K}KN\log_{2}N\right)$.
Note that from Theorem 4 and Remark 1, it follows that the proposed MSGT algorithm performs $\mathcal{O}\left(\frac{1}{L}\left(\frac{N}{\gamma K}\right)^{K}\right)$ times fewer computational operations compared to MAP.

### v5

@@ III-B Analytical Results
Let $G$ denote the number of non-defective items that were hidden and not detected in DND (i.e., the number of false positive items). We introduce the following upper-tail deviation probability for the number of false positives
$P_{e,u}^{(\mathrm{DND})}\;\triangleq\;P\!\left(G-\mathbb{E}[G]>g\,\big|\,K={\color[rgb]{0,0,0}k}\right),\quad\text{for}\quad g>0.$
Under the defective model given in Section II with $K_{\min}\leq K\leq K_{\max}$, Stage 2 enforces only an upper budget ${\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}$ on the support size (Algorithm 1). Accordingly, we choose $g$ to control the event that the number of possibly defective items after DND exceeds this budget, e.g.,
$g\;=\;{\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}\;-\;\mathbb{E}\!\left[\left|\mathcal{P}^{(S_{1,1})}\right|\mid K={\color[rgb]{0,0,0}k}\right]\;=\;{\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}\;-\;\Big({\color[rgb]{0,0,0}k+(N-k)}P_{e,a}^{(\mathrm{DND})}\Big),$
whenever this quantity is positive. Applying the one-sided Chebyshev’s inequality [27] to bound the concentration term gives
$P\!\big(G-\mathbb{E}[G]>g\,\big|\,K={\color[rgb]{0,0,0}k}\big)\;\leq\;\frac{\mathrm{Var}(G\mid K={\color[rgb]{0,0,0}k})}{\mathrm{Var}(G\mid K={\color[rgb]{0,0,0}k})+g^{2}}.$ (5)
Let $M_{0}$ denote the number of negative tests, $q_{0}\triangleq(1-p)^{{\color[rgb]{0,0,0}k}}$, and $n^{\prime}\triangleq N-{\color[rgb]{0,0,0}k}$. Each test is negative independently with probability $q_{0}$, hence $M_{0}\sim Bin(T,q_{0})$. Conditioned on $M_{0}=m_{0}$, a fixed non-defective item appears in none of the $m_{0}$ negative tests with probability $(1-p)^{m_{0}}$. Consequently, $G\,\big|\,M_{0}=m_{0},\ K={\color[rgb]{0,0,0}k}\;\sim\;Bin\big(n^{\prime},\ (1-p)^{m_{0}}\big)$ [3]. For $M\sim Bin(T,q)$ and any $a\in\mathbb{R}$,
$\mathbb{E}\!\left[a^{M}\right]\;=\;\big((1-q)+q\,a\big)^{T}.$ (6)
Using the law of total variance and $r\triangleq\left(1-p\right)^{M_{0}}$,
$\mathrm{Var}(G\mid K={\color[rgb]{0,0,0}k})$ $=\mathbb{E}\!\left[\mathrm{Var}(G\mid M_{0},K={\color[rgb]{0,0,0}k})\right]+\mathrm{Var}\!\left(\mathbb{E}[G\mid M_{0},K={\color[rgb]{0,0,0}k}]\right)$
$=\mathbb{E}\!\left[n^{\prime}r(1-r)\right]+\mathrm{Var}(n^{\prime}r)$
$=n^{\prime}\big(\mathbb{E}[r]-\mathbb{E}[r^{2}]\big)+n^{\prime 2}\big(\mathbb{E}[r^{2}]-(\mathbb{E}[r])^{2}\big).$ (7)
Using (6), we obtain
$\bar{r}_{1}=\mathbb{E}[r]=(1-pq_{0})^{T},\quad\text{and}\quad\bar{r}_{2}=\mathbb{E}[r^{2}]=\big(1-(2p-p^{2})q_{0}\big)^{T}.$
Substituting these expressions into the variance decomposition (7) yields the variance formula stated in (4). Finally, combining (5) with (4) establishes the upper-tail bound in (3), completing the proof.
@@
Consider noiseless nonadaptive group testing with $N$ items, a random number of defectives $K$ (conditioned on its realization), and $T$ tests, under a Bernoulli(p)(p) design with design parameter $p=\ln(2)/K_{\max}$. Fix any realization $k\in[K_{\min},K_{\max}]$ and consider performance conditioned on $\{K=k\}$. The number of defective items detected by the DD decoder is bounded by
$\mathbb{E}\big[\,|\mathcal{P}^{(\mathrm{DD})}|\,\big|\,K={\color[rgb]{0,0,0}k}\big]\;\geq\;{\color[rgb]{0,0,0}k}\!\left\{a\,T(1-q_{0})\,(1-p)^{N-{\color[rgb]{0,0,0}k}}-\frac{a^{2}}{2}\,T^{2}\right\}^{+},$ (8)
where $\{x\}^{+}\triangleq\max\{x,0\}$, all expectations are taken conditionally on the realized $K$, and $q_{0}\triangleq(1-p)^{{\color[rgb]{0,0,0}k}}$, $q_{1}\triangleq p(1-p)^{{\color[rgb]{0,0,0}k-1}}$, and $a\triangleq q_{1}/(1-q_{0})$.
Proof:.
The proof of Theorem 3 in given in Appendix B. ∎
Remark 2.
Under the Bernoulli design used in this work, the testing matrix is parameterized by a design sparsity range $K_{\min},K_{\max}]$, and the realized number of defectives $K$ is random and unknown to the decoder. Accordingly, the lower bound in (8), which is stated conditionally on a fixed realization of $K$, may evaluate to zero for some values of $K\in[K_{\min},K_{\max}]$ after applying the positive-part operator. As a result, the bound is not uniformly informative over the entire admissible sparsity range, but remains valid for all realizations of $K$. Since we always have the trivial bound $\mathbb{E}[\,|\mathcal{P}^{(\mathrm{DD})}|\,]\geq 0$, this behavior is consistent with correctness and does not indicate a failure of the analysis. Moreover, the proposed 2SDGT algorithm does not rely on the magnitude of $|\mathcal{P}^{(\mathrm{DD})}|$ for correctness or feasibility. In particular, Stage 2 applies a deterministic filtering rule that retains only candidate supports whose cardinality lies in the range $K_{\min}\leq|\mathcal{V}^{(l)}|\leq\lfloor\gamma K_{\max}\rfloor$, independently of the quantitative DD bound. For this reason, the complexity analysis of 2SDGT in Theorem 4 relies only on the elementary inequalities $0\leq|\mathcal{P}^{(\mathrm{DD})}|\leq|\mathcal{P}^{(S_{1,1})}|\leq N$, rather than on the quantitative lower bound (8).
The proposition below states a prior-free counting bound relating the Stage 2 candidate-set budget and the best achievable uniformly averaged probability of exact support recovery at a fixed sparsity level $k$, where the averaging is uniform over all supports of cardinality $k$ (see Remark 3). It is used to motivate how the Stage 2 parameter $\gamma$ should scale with $T$ and $(N,K_{\max})$. Although the 2SDGT and MAP decoders explicitly depend on the prior $P(\mathcal{K})$, this proposition is a prior-free counting bound that applies to any decoding rule. In particular, the bound remains valid when MAP is used with an arbitrary correlated prior, since it constrains only the number of candidate supports consistent with a given outcome.
Proposition 1.
Fix any $k\in\{K_{\min},\ldots,K_{\max}\}$. Let $\eta_{k}\triangleq\left({1}/{\binom{N}{k}}\right)\sum_{\mathcal{K}:\,|\mathcal{K}|=k}P\!\left(\hat{\mathcal{K}}=\mathcal{K}\right)$ denote the uniformly averaged (over all $k$-subsets) probability of exact support recovery at sparsity level $k$. Consider any decoding procedure that, given $Y$, outputs a candidate item set $\mathcal{L}(Y)\subseteq[N]$ satisfying $|\mathcal{L}(Y)|\leq{\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}$. Then, achieving $\eta_{k}\geq\bar{\eta}_{k}$ requires
$T\;\geq\;\log_{2}\bar{\eta}_{k}+\log_{2}\binom{N}{k}-\log_{2}\binom{{\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}}{k}$. (9)
@@
Fix an arbitrary $k\in\{K_{\min},\ldots,K_{\max}\}$. We follow a standard counting argument in the spirit of list-decoding converses (cf. [54]), adapted here to decoders that output a bounded-size item set. For each outcome vector $Y\in\{0,1\}^{T}$, define
$N(Y)\triangleq\sum_{\mathcal{K}:\,|\mathcal{K}|=k}\mathbb{1}\{\mathcal{K}\subseteq\mathcal{L}(Y)\},$
namely, the number of $k$-supports fully contained in the candidate set $\mathcal{L}(Y)$.
Since $|\mathcal{L}(Y)|$ is integer, we have $|\mathcal{L}(Y)|\leq\lfloor\gamma K_{\max}\rfloor$, and hence
$N(Y)=\binom{|\mathcal{L}(Y)|}{k}\leq\binom{\lfloor\gamma K_{\max}\rfloor}{k}$.
Moreover, for any true support $\mathcal{K}$ with $|\mathcal{K}|=k$, the event $\{\hat{\mathcal{K}}=\mathcal{K}\}$ necessarily implies that all defective items must survive Stage 2, and hence $\mathcal{K}\subseteq\mathcal{L}(Y)$. Averaging uniformly over all such supports and summing over all outcomes yields
$\eta_{k}\leq\frac{1}{\binom{N}{k}}\sum_{Y}N(Y),$
where the inequality follows since $\mathbb{1}\{\hat{\mathcal{K}}=\mathcal{K}\}\leq\mathbb{1}\{\mathcal{K}\subseteq\mathcal{L}(Y)\}$ and $\Pr(Y\mid\mathcal{K})\leq 1$.
Using the bound on $N(Y)$ and the fact that there are $2^{T}$ possible outcomes, we obtain
$\eta_{k}\leq\frac{2^{T}}{\binom{N}{k}}\binom{\lfloor\gamma K_{\max}\rfloor}{k}$. (10)
Rearranging (10) completes the proof.
Remark 3.
The success probability $\eta_{k}$ in Proposition 1 is defined as the probability of exact support recovery, uniformly averaged over all supports of cardinality $k$. This metric is not the prior-dependent success probability of the MAP/LVA decoder under the correlated prior considered in this work. It is introduced solely to enable a prior-free counting argument. Importantly, this proposition does not characterize the performance of MAP decoding under the correlated prior. Rather, it provides a necessary combinatorial limitation that applies to any decoding rule whose Stage 2 output is restricted to a candidate item set of size at most ${\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}$. Since this limitation is independent of the prior distribution over supports, it remains valid when MAP is employed with an arbitrary correlated prior. The bound is therefore used solely to motivate the scaling of $\gamma$ and $T$, and not as a direct performance guarantee.
One of the key features of the proposed 2SDGT algorithm is its low and feasible complexity in practical regimes compared to ML or MAP-based GT decoders. Both ML and MAP involve exhaustive searches, resulting in a complexity of $\mathcal{O}\left(\binom{N}{K}KN\log_{2}N\right)$ operations [8]. The theorem below and the subsequent Remark characterize the computational complexity of 2SDGT.
Theorem 4.
Consider a group test for a population of $N$ items, with an unknown number of defectives $K$ satisfying $K_{\min}\leq K\leq K_{\max}$, and a Bernoulli testing matrix designed using $K_{\max}$. The computational complexity of the 2SDGT algorithm is bounded by $\color[rgb]{0,0,0}\mathcal{O}\bigl(L\gamma^{K_{\max}}K_{\max}N\log_{2}N\bigr)$} operations.
@@
We begin by analyzing the complexity of each step of the proposed 2SDGT solution given in Algorithm 1, and finally sum everything up to determine the total complexity.
@@
The complexity of DND is $\color[rgb]{0,0,0}\mathcal{O}\bigl(K_{\max}N\log_{2}{N}\bigr)$}, as analyzed in [19, Remark 6]. Then, for each positive entry of the test result vector $Y^{T}$, the DD algorithm counts the number of possibly defective items that participate in the corresponding pool test. This can be implemented with at most ${\color[rgb]{0,0,0}K_{\max}}\bigl|\mathcal{P}^{(S_{1,1})}\bigr|\log_{2}{N}$ operations, which we bound by the DND complexity, i.e., by $\color[rgb]{0,0,0}\mathcal{O}\bigl(K_{\max}N\log_{2}N\bigr)$}. Hence, the overall complexity of Stage 1 is $\color[rgb]{0,0,0}\mathcal{O}\bigl(K_{\max}N\log_{2}N\bigr)$}.
Parallel LVA requires $L$ times more computations than the Viterbi Algorithm (VA) [56]. The VA calculates all the possible transition probabilities for each step in the sequence. In GT, this sequence is the items’ ordering, and with the proposed algorithm it is enough to consider only the $\bigl|\mathcal{P}^{(S_{1,2})}\bigr|$ items as the sequence length.
The possible states are “non-defective” and “defective”, so there are four transitions in each step of the GT trellis. More generally, the algorithm can be implemented to leverage additional memory and decide the state of each item based on the preceding $\tau$ items. Consequently, LVA takes at most
$2^{2\tau}L\bigl|\mathcal{P}^{(S_{1,2})}\bigr|$
elementary operations. For the purpose of an upper bound on the complexity, it suffices to note that for every realization
$0\,\leq\,\bigl|\mathcal{P}^{(DD)}\bigr|\,\leq\,\bigl|\mathcal{P}^{(S_{1,1})}\bigr|\,\leq\,N,$
and hence
$\bigl|\mathcal{P}^{(S_{1,2})}\bigr|\;=\;\bigl|\mathcal{P}^{(S_{1,1})}\bigr|-\bigl|\mathcal{P}^{(DD)}\bigr|\;\leq\;\bigl|\mathcal{P}^{(S_{1,1})}\bigr|\;\leq\;N.$
Therefore, we have
$2^{2\tau}L\bigl|\mathcal{P}^{(S_{1,2})}\bigr|\;\leq\;2^{2\tau}LN\;=\;\mathcal{O}(LN),$
so the complexity of the LVA step in 2SDGT is $\mathcal{O}(LN)$.
@@
In the next step, we filter the LVA results. We sum each sequence $\mathbf{Z}_{\ell}$ with complexity $\mathcal{O}(N)$ per sequence. If the sum lies in the range $K_{\min},{\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}]$, we extract all combinations of sizes $k-\bigl|\mathcal{P}^{(DD)}\bigr|$ for all $k\in[K_{\min},K_{\max}]$. Hence, this step requires at most
$\color[rgb]{0,0,0}\mathcal{O}\!\left(L\left(N+\sum_{k=K_{\min}}^{K_{\max}}\binom{{\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}}{k-\left|\mathcal{P}^{(DD)}\right|}\right)\right)$.}
This term will be dominated by the subsequent MAP step.
Finally, in the MAP step of Stage 2, the algorithm considers at most $L$ sequences. In each sequence, there are at most ${\color[rgb]{0,0,0}\lfloor}{\color[rgb]{0,0,0}\gamma K_{\max}}{\color[rgb]{0,0,0}\rfloor}$ possibly defective items, and we examine all combinations of sizes $k-\bigl|\mathcal{P}^{(DD)}\bigr|$ for $k\in[K_{\min},K_{\max}]$. Thus, the number of candidate supports per sequence is at most
$\color[rgb]{0,0,0}\sum_{k=K_{\min}}^{K_{\max}}\binom{{\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}}{k-\left|\mathcal{P}^{(DD)}\right|}\;\leq\;\sum_{k=K_{\min}}^{K_{\max}}\binom{{\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}}{k}$.}
For each candidate support, we apply the group test metric, whose computation can be implemented with $\color[rgb]{0,0,0}\mathcal{O}\bigl(K_{\max}N\log_{2}N\bigr)$} operations. Therefore, the complexity of the MAP stage is bounded by
$\color[rgb]{0,0,0}L\sum_{k=K_{\min}}^{K_{\max}}\binom{{\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}}{k}K_{\max}N\log_{2}N.$}
Using the standard bounds
$\Big(\tfrac{{\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}}{K_{\max}}\Big)^{K_{\max}}\;\leq\;\binom{{\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}}{K_{\max}}\;\leq\;\Big(\tfrac{e{\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}}{K_{\max}}\Big)^{K_{\max}},$
we obtain
$L\sum_{k=K_{\min}}^{K_{\max}}\binom{{\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}}{k}K_{\max}N\log_{2}N\;=\;\mathcal{O}\bigl(L\gamma^{K_{\max}}K_{\max}N\log_{2}N\bigr).$
Taking all stages together, the overall complexity of 2SDGT is
$\mathcal{O}\Big(K_{\max}N\log_{2}N+LN+L\!\sum_{k=K_{\min}}^{K_{\max}}\binom{{\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}}{k-\left|\mathcal{P}^{(DD)}\right|}+L\gamma^{K_{\max}}K_{\max}N\log_{2}N\Big).$
As $N$ grows, this sum is dominated by the MAP term, and we conclude that the complexity of the 2SDGT algorithm is bounded by
$\color[rgb]{0,0,0}\mathcal{O}\bigl(L\gamma^{K_{\max}}K_{\max}N\log_{2}N\bigr)$},
which completes the proof. ∎
@@
Under the interval-sparsity formulation $K_{\min}\leq K\leq K_{\max}$, skipping the LVA step results in an exhaustive MAP search over all defective sets with cardinality in $K_{\min},K_{\max}]$, whose complexity scales as
$\mathcal{O}\!\left(\sum_{k=K_{\min}}^{K_{\max}}\binom{N}{k}\,K_{\max}N\log_{2}N\right),$
which remains combinatorial in the effective sparsity level.
Note that from Theorem 4 and Remark 4, it follows that the proposed 2SDGT algorithm performs
$\mathcal{O}\!\left(({1}/{L})\left({N}/{{\color[rgb]{0,0,0}\lfloor}\gamma K_{\max}{\color[rgb]{0,0,0}\rfloor}}\right)^{K_{\max}}\right),$
times fewer computational operations compared to exhaustive MAP search over $K_{\min},K_{\max}]$. The MAP baseline reported herein corresponds to MAP decoding performed over the reduced candidate set obtained after the Stage 1 pruning step, rather than exhaustive MAP decoding over all $\binom{N}{K}$ supports. This choice reflects the fact that exhaustive MAP decoding is computationally infeasible for the problem dimensions considered and would not provide a meaningful practical benchmark. The reported complexity, therefore, accounts only for the enumeration and likelihood evaluation over the retained candidate supports, enabling a fair comparison with the proposed 2SDGT algorithm.
The Bernoulli test-design parameter $p=\ln(2)/K_{\max}$ and the subsequent complexity and performance guarantees in this section are derived under a design sparsity interval $(K_{\min},K_{\max})$, rather than assuming knowledge of the realized sparsity $K$. This distinction is crucial: although the realized number of defectives enters the analysis of Stage 1 through concentration bounds, Stage 2 and the associated list-decoding and counting guarantees depend only on the design interval. Specifically, the LVA filtering rule in Algorithm 1 retains trajectories whose support size satisfies $K_{\min}\leq|V^{(\ell)}|\leq\lfloor\gamma K_{\max}\rfloor$, independently of the true realization of $K$. As a result, a moderate mismatch between the realized sparsity and the design upper bound affects performance only through the tail probability $\Pr(K<K_{\min})+\Pr(K>K_{\max})$, which is explicitly separated in the overall error bound in (1). Importantly, the counting argument in Proposition 1 and the complexity bound in Theorem 4 remain valid regardless of the exact value of $K$, as they depend solely on the candidate-set budget $\lfloor\gamma K_{\max}\rfloor$ and not on the decoder’s knowledge of the realized sparsity. This formulation enables 2SDGT to operate reliably under random or imperfectly known sparsity levels without redesigning the decoding pipeline, while preserving provable complexity control.